OpenStudy (anonymous):

Find a point on the y-axis that is equidistant from the points (4, −4) and (1, 1).

OpenStudy (anonymous):

The point will be (0, y). x is zero because it is on the y-axis. To proceed from here, we use Pythagoras' Theorem. \[d^{2}=(4-0)^{2}+(-4-y)^{2}\] \[d^{2}=(1-0)^{2}+(1-y)^{2}\] These expressions are equal so: \[y ^{2}+8y+32=y^{2}-2y+2\] \[y=-3\] Your point is \[P(0,-3)\]

OpenStudy (anonymous):

thanku